# Scalable LLM Math Reasoning Acceleration with Low-rank Distillation

### v1

@@ Abstract
Due to long generations, large language model (LLM) math reasoning demands significant computational resources and time. While many existing efficient inference methods have been developed with excellent performance preservation on language tasks, they often severely degrade math performance. In this paper, we propose Caprese, a resource-efficient distillation method to recover lost capabilities from deploying efficient inference methods, focused primarily in feedforward blocks. With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the math capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs. Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>11% reduction to generate 2048 tokens with Qwen 2.5 14B) while encouraging response brevity.
@@ 1 Introduction
With the increasing capabilities of large language models (LLMs) [VSP+17], evaluations are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving. These tasks tend to demand long generation which drives up latency, making efficiency a dire issue. Fortunately, many efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization. However, for math reasoning tasks, many efficient inference algorithms begin to break down, decimating performance, despite their robustness in language settings. Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and math capabilities.
@@
One of the main differences between math reasoning and many language tasks is the generation length. Reasoning typically involves long generation due to improved performance from chain-of-thought (CoT) [WWS+22], which often surpasses the length of the input query. However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time. For instance, deploying CATS [LLZ+24], a sparse thresholding method, on Gemma 2 2B [TRP+24] has little impact on language generation performance, but knocks GSM8K [CKB+21] accuracy from 51.02% to 34.42% (Table 1). Similarly, for thinking models: applying GRIFFIN [DCC24], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [GYZ+25] drops MATH-500 [LKB+23] accuracy from 79.40% to 42.00% (Table 2). Moreover, since generation is much more demanding computationally than prefill (e.g., >150$\times$ slower to generate than prefill 1K tokens on one GPU with Llama 3.1 8B), there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling. Thus, this poses two key inference challenges with math reasoning:
@@
    Instability: Long CoTs can be sensitive to mistakes. Even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24]. Hence, efficient inference algorithms should respect the delicacy of CoT.
@@
    Inefficiency: Due to the reliance on CoT, which tend to lead to long generation, the already computationally expensive LLM autoregressive decoding process is accentuated. This problem is further exaggerated as CoTs scale in length in thinking models [GYZ+25] and as the number of repeated generations scale to elicit higher quality answers [BJE+24, WSL+24, SLXK24].
Figure 1: A full FF block will contain the complete performance of the original model without any benefit to efficiency. Sparse FF algorithms can be very efficient by using subsets of the FF block but harm math performance. Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for highly performative inference in language and math settings while being efficient. Layers are drawn as trapezoids to highlight the relative size of inputs and outputs.
An ideal method should be efficient, performance preserving, and easy to integrate. Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible. (We choose to focus on FF blocks since, contributing around 2/3 of an LLM’s parameters and over 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods. Using an oracle top-$k$ filter on FF nonlinearity output magnitudes, we observe huge reductions in error with a low-rank approximation to the FF output residuals (Figure 2). Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively mundane. This observation motivates us to estimate the residual from sparse FF methods with low-rank layers.
Figure 2: Average relative FF output error of generated tokens with varying top-$k$ densities and low-rank approximations. The density is the fraction of non-zero intermediate FF features maintained by top-$k$. Samples consist of 16 random MATH prompts and generations by the original model. A relatively small low-rank approximation to the top-$k$ residual dramatically reduces error much more effectively than just increasing density.
We introduce Caprese [^1] Our method’s components are modular with easy preparation like a Caprese salad. (CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1). Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome both challenges of math reasoning and makes significant progress towards an ideal efficient method:
  1. 1.
    Performance Enhancement: Demonstrated across a variety of models (e.g., instruct and thinking models), Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks. Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling. For example, scaling generations with Caprese on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
  2. 2.
    Efficiency: Due to the ability to combine our low-rank layers with existing FF layers, Caprese reduces generation latency. For instance, Caprese reduces latency by 11.7% when generating 2048 tokens using DeepSeek-R1-Distill-Qwen 14B with GRIFFIN.
@@
While Caprese is not restricted to math tasks, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities. Future work can investigate its benefit to other domains.
The remainder of this paper is organized as follows. Section 2 covers background on related works and describes GRIFFIN and CATS, two efficient sparse FF algorithms that serve as baselines. Section 3 details Caprese’s architecture and distillation procedure. Then, we showcase the strong performance of Caprese in Section 4: instruct LLM inference (Section 4.1), scaling generation outputs (Section 4.2), and scaling generation length with thinking LLMs (Section 4.3). Finally, we quantify latency improvements in Section 4.4.
@@ 2.1 Related Works
To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23]. Pruning sets a parameter subset to zero to enforce static sparsity [LDS89, SLBK23, FA23, MFW23]. Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [JJNH91, ZLL+21, DDZ+24, LWD+23]. Similar to MoEs, another line of work aims at exploiting contextual sparsity to dynamically select FF neurons for each input without training. GRIFFIN [DCC24] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation. CATS [LLZ+24] uses hard thresholding to skip computation in parts of the FF block with a custom kernel. More details of GRIFFIN and CATS can be found in Section 2.2.
@@
To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [KMH+20, HBM+22]. More recently, there is an increasing effort towards test-time scaling to boost performance on difficult tasks like math. Two ways of test-time scaling that have seen great success. First, for a single prompt, multiple responses can be sampled from the LLM [BJE+24, SLXK24, WSL+24, MSE24, SHZ+24]. This increases the probability that a desired response lies in the pool of responses. While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one. Second, CoTs can be lengthened for each response [GYZ+25]. By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
@@ 2.2 Training-free & Adaptively Sparse FF Methods
GRIFFIN [DCC24] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns. GRIFFIN calculates $[\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $[\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$. The result $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ gives a metric to perform top-$k$ selection across corresponding columns of $\bm{W}_{g}$ and $\bm{W}_{1}$, and rows of $\bm{W}_{2}$ to produce $\widehat{\bm{W}}_{g},\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{k\times D}$. Then, for the generation phase, the following FF block is used for input $\bm{x}\in\mathbb{R}^{D}$:
@@ 3.1 Layer-wise Distillation
for input set $\mathcal{X}$, $\bm{L}\in\mathbb{R}^{D\times r}$, and $\bm{R}\in\mathbb{R}^{r\times D}$ where $r\ll D_{\text{FF}}$. The optimal solution can be computed analytically since this is a reduced rank regression problem, but the size of $|\mathcal{X}|$ and $D$ may make it prohibitively expensive. Therefore, we opt to learn $\bm{L}$ and $\bm{R}$ independently for every FF block (i.e., previous FF blocks are assumed to be from the original model), allowing for parallel layer-wise training. This takes inspiration from LESS [DYZ+24] which uses layer-wise training on attention residuals for key-value cache compression. Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen.
Even if the layer-wise MSE is minimized, this does not guarantee the final outputs closely resemble the original model due to error accumulation. Hence, we also distill end-to-end (E2E) to further improve the performance.
@@ 3.2 End-to-end Distillation
where M and $\text{M}_{\text{student}}$ are the original LLM and efficient LLM with distillation layers, respectively, excluding the final linear head. $L$ is the number of layers in the the model. All original weights are frozen, so the only tunable parameters are the $\bm{L}$ and $\bm{R}$ of each FF block.
@@ 3.4 Training Details
We set the inner dimension of our low-rank layer to $r=256$ (to see the effect of different $r$, see Appendix A). In comparison to the enormous inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B).
We use a 20K subset of a common synthetic math training set for training. For a fair comparison, the same subset is used for both layer-wise and E2E distillation for every model. Each training sample is prepended with a CoT instruction: “Please reason step by step.” At test time, the actual instructions may be vastly different. Layer-wise and E2E training consists of 20 epochs and 3 epochs, respectively. Training and inference are done in BF16. Hyperparameters are listed in Appendix E.
@@
Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). FF sparsity is set to 50%, and $r=256$.
@@ 4 Experiments
We showcase the effectiveness of Caprese at recovering much, if not all, of the math performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks. In the following section (Section 4.1), we observe better math performance for instruct LLMs without losing their generalizability. Then, we study Caprese’s benefit on two axes of test-time scaling: number of generations (Section 4.2) and generation length with thinking models (Section 4.3). Finally, we highlight Caprese’s latency improvements in Section 4.4. When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying sparse method and similarly for GRIFFIN. Otherwise, we use “Caprese” for brevity. Unless specified, the FF intermediate feature sparsity is set at 50%. NOTE: For a more meaningful baseline, we only apply GRIFFIN to the first half of the model for all experiments since we observed much steeper drops in math performance if used for all layers. Regardless, Caprese still improves the performance of GRIFFIN on all layers. CATS is applied to all layers.

### v2

@@ Abstract
Due to long generations, large language model (LLM) math reasoning demands significant computational resources and time. While many existing efficient inference methods have been developed with excellent performance preservation on language tasks, they often severely degrade math performance. In this paper, we propose Caprese, a resource-efficient distillation method to recover lost capabilities from deploying efficient inference methods, focused primarily in feedforward blocks. With original weights unperturbed, roughly 1% of additional parameters, and only 20K synthetic training samples, we are able to recover much if not all of the math capabilities lost from efficient inference for thinking LLMs and without harm to language tasks for instruct LLMs. Moreover, Caprese slashes the number of active parameters ($\sim$2B cut for Gemma 2 9B and Llama 3.1 8B) and integrates cleanly into existing model layers to reduce latency (>16% time-to-next-token reduction) while encouraging response brevity (up to 8.5% fewer tokens).
@@ 1 Introduction
With the increasing capabilities of large language models (LLMs) [VSP+17], evaluations are also becoming increasingly sophisticated, typically involving multi-step reasoning such as in math problem solving. These tasks tend to demand long generation which drives up latency, making efficiency a dire issue. Fortunately, many sparsity-based efficient LLM inference algorithms have shown great promise, slashing expensive computational bottlenecks with little damage to the original performance on a variety of language-based tasks like reading comprehension and summarization. However, for math reasoning tasks, many of these algorithms begin to break down, decimating performance, despite their robustness in language settings. Thus, there is a need to design efficient inference algorithms that simultaneously maintain language and math capabilities.
@@
One of the main differences between math reasoning and many language tasks is the generation length. Reasoning typically involves long generations from chain-of-thoughts (CoTs) [WWS+22], which often far surpass the length of the input query. However, CoT performance falls apart with efficient algorithms that introduce approximation errors which build up over time. For instance, deploying CATS [LLZ+24], a sparse thresholding method, on Gemma 2 2B [TRP+24] has little impact on language generation performance, but knocks GSM8K [CKB+21] accuracy from 51.02% to 34.42% (Table 1). Similarly, for thinking models: applying GRIFFIN [DCC24], an adaptive structured pruning method, on the first half of DeepSeek-R1-Distill-Qwen 1.5B [GYZ+25] drops MATH-500 [LKB+23] accuracy from 79.40% to 42.00% (Table 2). Moreover, since generation is much more computationally demanding than prefill, there is a dire need to make long reasoning CoTs efficient, which is made more apparent with the rise of test-time scaling. Thus, this poses two key inference challenges with math reasoning. First, even single token mistakes can sometimes drive the generation trajectory off course, leading to an incorrect response [ZCX+24]. Second, the already computationally expensive LLM autoregressive decoding process is accentuated with generating CoTs. This problem is further exaggerated as CoTs scale in length in thinking models [GYZ+25] and as the number of repeated generations scale to elicit higher quality answers [BJE+24, WSL+24, SLXK24]. An ideal method should be efficient, performance preserving, and easy to integrate.
Figure 1: A full FF block maximizes accuracy without any benefit to efficiency. Sparse FF algorithms can be very efficient by using subsets of the FF block but harm math performance. Our method, Caprese, uses a sparse FF algorithm and a small distilled low-rank linear layer, which can be merged with existing FF weights, for performative inference in language and math settings while being efficient. Layers are drawn as trapezoids to highlight the relative size of inputs and outputs.
Thankfully, low-rank structure in the feedforward (FF) output features can help make this possible. (We choose to focus on FF blocks since they contribute around 2/3 of an LLM’s parameters and about 50% of the generation latency.) Because of the success of works that exploit contextual sparsity in FF blocks on language tasks for efficiency, like CATS and GRIFFIN, we peek into the residuals of FF sparsity-based efficient methods. Using an oracle top-$k$ filter on FF nonlinearity output magnitudes, we observe huge reductions in error with a low-rank approximation to the FF output residuals (Figure 2). Since FF intermediate feature sizes can be on the order of $10^{5}$, adding 256 for a low-rank approximation is comparatively tiny. This observation motivates us to estimate the residual from sparse FF methods with low-rank layers.
Figure 2: Average relative FF output error of generated tokens with varying top-$k$ densities and low-rank approximations. The density is the fraction of non-zero intermediate FF features maintained by top-$k$. Samples consist of 16 random MATH generations by the original model. A relatively small low-rank approximation to the top-$k$ residual reduces error more effectively than increasing density.
We introduce Caprese (CAPability REcovery with Scalable Efficiency) to learn FF residuals from any sparse FF algorithm using small low-rank linear layers for improved performance (Figure 1). While Caprese works for any task, we focus on math in this paper, as current efficient sparse methods can obliterate an LLM’s math capabilities. Through distillation of math knowledge into these low-rank layers, Caprese is able to overcome the aforementioned challenges of math reasoning and makes significant progress towards an ideal efficient method:
@@
    Performance Enhancement: Demonstrated across a variety of instruct and thinking models, Caprese recovers much if not all of the math performance lost from deploying a sparse FF algorithm without harming language tasks. Also, Caprese maintains its performance benefit when applying different techniques of test-time scaling.
@@
    Efficiency: Due to its parallelizability with existing FF layers, Caprese adds negligible overhead, preserving latency reductions of the underlying method.
@@
Our extensive experiments on Caprese indicate strong performance on various tasks and models. For example, DeepSeek-R1-Distill-Qwen 7B with a sparse method drops AMC 2023 accuracy from 75.00% to 62.50%, but Caprese is able to bring it to 78.13%, beyond the original accuracy. Furthermore, scaling generation quantity with Caprese on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% fewer active parameters. We also show Caprese reduces latency by 16% using the Qwen 2.5 14B architecture.
Paper Organization.
Section 2 covers background on related works and describes GRIFFIN and CATS, two efficient sparse FF algorithms that serve as baselines. Section 3 details Caprese’s architecture and distillation procedure. Then, we showcase the strong performance of Caprese in Section 4: instruct LLM inference (Section 4.1), scaling generation outputs (Section 4.2), and scaling generation length with thinking LLMs (Section 4.3). Finally, we quantify efficiency improvements in Section 4.4.
@@ 2.1 Related Works
To improve the efficiency of FF blocks in LLMs, various methods leverage existing sparse structures in FF features [GSBL20, DLBZ22, LYB+22, DCC23, LWD+23]. Pruning sets a parameter subset to zero to enforce static sparsity [LDS89, SLBK23, FA23, MFW23]. Mixtures of experts (MoEs) adaptively select predefined parameter subsets in FF blocks, though they tend to require significant fine-tuning or training from scratch [JJNH91, ZLL+21, DDZ+24, LWD+23]. Similar to MoEs, another line of work aims at exploiting contextual sparsity to dynamically select FF neurons for each input without training. GRIFFIN [DCC24] is a calibration-free method that uses FF activations from the prefill phase to adaptively prune FF neurons for generation. CATS [LLZ+24] uses hard thresholding to skip computation in parts of the FF block with a custom kernel.
@@
To enhance LLM performance, the priority has historically been to scale training with more data and bigger models [KMH+20, HBM+22]. More recently, there is an increasing effort towards test-time scaling to boost performance on difficult tasks like math. Two ways of test-time scaling that have seen great success. First, for a single prompt, multiple responses can be sampled from the LLM [BJE+24, SLXK24, WSL+24, MSE24, SHZ+24]. This increases the probability that a desired response lies in the pool of responses. While this method of scaling is highly parallelizable, it also relies on a verifier model to select the best one. Second, CoTs can be lengthened for each response [GYZ+25, MYS+25, YLY+25]. By extrapolating the success of CoTs to extreme lengths (e.g., DeepSeek-R1 generates up to 32K tokens per response), the accuracy of the final answer improves significantly, but this method has low parallelizability due to the autoregressive nature of generation.
@@ 2.2 Training-free & Adaptively Sparse FF Methods
GRIFFIN [DCC24] adaptively prunes columns and rows in the FF block weights, using FF activation statistics from the prefill phase, namely the flocking patterns. GRIFFIN calculates $\overline{\bm{Z}}]_{i}=|[\bm{Z}]_{i}|/\|[\bm{Z}]_{i}\|_{2}$ for each token index $i$, followed by an aggregation across the FF feature axis: $\bm{s}]_{j}=\|[\overline{\bm{Z}}]_{\cdot,j}\|_{2}$. The result $\bm{s}\in\mathbb{R}^{D_{\text{FF}}}$ gives a metric to perform top-$k$ selection across corresponding columns of $\bm{W}_{g}$ and $\bm{W}_{1}$, and rows of $\bm{W}_{2}$ to produce $\widehat{\bm{W}}_{g},\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$, and $\widehat{\bm{W}}_{2}\in\mathbb{R}^{k\times D}$. Then, for the generation phase, the following FF block is used for input $\bm{x}\in\mathbb{R}^{D}$:
@@ 3.1 Layer-wise Distillation
for input set $\mathcal{X}$, $\bm{L}\in\mathbb{R}^{D\times r}$, and $\bm{R}\in\mathbb{R}^{r\times D}$ where $r\ll D_{\text{FF}}$. The optimal solution can be computed analytically since this is a reduced rank regression problem, but the size of $|\mathcal{X}|$ and $D$ may make it prohibitively expensive. Therefore, we opt to learn $\bm{L}$ and $\bm{R}$ independently for every FF block (i.e., previous FF blocks are assumed to be from the original model), allowing for parallel layer-wise training. This takes inspiration from LESS [DYZ+24] which uses layer-wise training on attention residuals for key-value cache compression. Each $\bm{R}$ is initialized as a zero matrix since the efficient approximation is assumed to be of good quality, and original model weights are frozen. We also distill end-to-end (E2E) to further improve the performance.
@@ 3.2 End-to-end Distillation
where M and $\text{M}_{\text{student}}$ are the original LLM and efficient LLM with distillation layers, respectively, excluding the final linear head. $L$ is the number of layers in the model. All original weights are frozen, so the only tunable parameters are the $\bm{L}$ and $\bm{R}$ of each FF block.
@@ 3.4 Training Details
We set the inner dimension of our low-rank layer to $r=256$ (to see the effect of different $r$, see Appendix A). In comparison to the large inner dimension of FF layers (e.g., $D_{\text{FF}}=14336$ for Llama 3.1 8B and Gemma 2 9B), our choice of $r$ is relatively miniscule, adding only $\sim 1\%$ new parameters for all tested models (Appendix B). We use a 20K subset of a common synthetic math training set for training. For a fair comparison, the same subset is used for both layer-wise and E2E distillation for every model. Each training sample is prepended with a CoT instruction: “Please reason step by step.” At test time, the actual instructions may be vastly different. Layer-wise and E2E training consists of 20 epochs and 3 epochs, respectively. Training and inference are done in BF16. Hyperparameters are listed in Appendix E.
3.5 Comparison with LoRA
Our method derives inspiration from and shares a slight connection with low-rank adaptation (LoRA), a widely used method for efficient fine tuning with the addition of low-rank parameters [HSW+22], but remains distinct since they target different inefficiencies. To illustrate, recall a sparse algorithm constructs $\widehat{\bm{W}}_{1}\in\mathbb{R}^{D\times k}$ by selecting columns from $\bm{W}_{1}$. Alternatively, construct $\widetilde{\bm{W}}_{1}\in\mathbb{R}^{D\times D_{\text{FF}}}$ by setting unwanted columns to zero, so nonzero columns in $\widetilde{\bm{W}}_{1}$ match columns in $\widehat{\bm{W}}_{1}$ and vice versa. Doing the same to columns of $\bm{W}_{g}$ and rows of $\bm{W}_{2}$, LoRA learns $\bm{A}_{1},\bm{A}_{g},\bm{B}_{2}^{\top}\in\mathbb{R}^{D\times r}$ and $\bm{B}_{1},\bm{B}_{g},\bm{A}_{2}^{\top}\in\mathbb{R}^{r\times D_{\text{FF}}}$ to form a new FF block with LoRA weights:
$\widetilde{\text{FF}}(\bm{x})=\left(\sigma\big(\bm{x}(\widetilde{\bm{W}}_{g}+\bm{A}_{g}\bm{B}_{g})\big)\odot\big(\bm{x}(\widetilde{\bm{W}}_{1}+\bm{A}_{1}\bm{B}_{1})\big)\right)(\widetilde{\bm{W}}_{2}+\bm{A}_{2}\bm{B}_{2}).$ (11)
From this, we see LoRA is not designed to be applied to sparse efficient inference algorithms. First, this is because in its compressed form, LoRA requires two sequential operations per linear layer. In contrast, Caprese parameters can be appended to existing ones, so additional computation is all parallel to the original operations (Section 3.3). The real latency impact of this difference is quantified later in Table 4. Caprese’s architecture also allows efficient training, but that is merely a byproduct of our design, not its purpose. Second, for a method like GRIFFIN which reduces the intermediate FF feature size from $D_{\text{FF}}$ to $k$, adding LoRA will negate this benefit while Caprese will preserve the reduced feature size. Due to this difference, we de-prioritize comparisons with LoRA in the main text, but include comparisons with LoRA in Appendix C for interested readers.
@@ 3.5 Comparison with LoRA
Table 1: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER). $F$ sparsity is set to 50%, and $r=256$.
@@ 4 Experiments
We showcase the effectiveness of Caprese at recovering much, if not all, of the math performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks. In the following section (Section 4.1), we observe better math performance for instruct LLMs without losing their generalizability. Then, we study Caprese’s benefit on two axes of test-time scaling: number of generations (Section 4.2) and generation length with thinking models (Section 4.3). Finally, we highlight Caprese’s latency and length improvements in Section 4.4. When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying sparse method and similarly for GRIFFIN. Otherwise, we use “Caprese” for brevity. Unless specified, the FF intermediate feature sparsity is set at 50%. NOTE: For a more meaningful baseline, we only apply GRIFFIN to the first half of the model for all experiments since we observed much steeper drops in math performance if used for all layers. CATS is applied to all layers.